A firm's average total cost is $80, its average variable cost is $75, and its output is 50 units. Its total fixed cost is:


1
Expert's answer
2021-03-08T09:19:16-0500

By the definition of the average total cost, we have:


"ATC=AFC+AVC,""ATC=\dfrac{TFC}{Q}+AVC,""TFC=(ATC-AVC)Q,""TFC=(\$80-\$75)\cdot50=\$250."